# Active Advantage-Aligned Online Reinforcement Learning with Offline Data

### v2

@@ 4.1 Confidence-aware Active Advantage-Aligned Sampling Strategy
Estimating the likelihoods $d^{\text{off}}\left({s,a}\right)$ and $d^{\text{on}}\left({s,a}\right)$ poses a challenge, as they could represent stationary distributions from mixture of complex policy. To address this issue, we employ a method studied by Lee et al. [29], Sinha et al. [58] for density ratio estimation that does not rely on likelihoods. This method approximates ${w}\left({s,a}\right)$ by training a neural network ${w}_{\psi_{i}}\left({s,a}\right)$, which is parameterized by $\psi_{i},i\in[N_{e}]$, where $N_{e}$ is the number of density networks in the ensemble. The training exclusively uses samples from $\mathcal{B}^{\text{off}}$ and $\mathcal{B}^{\text{on}}$. We use variational representation of f-divergences [44]. Consider $P$ and $Q$ as probability measures on a measurable space $\mathcal{X}$, with $P$ being absolutely continuous $w.r.t$ $Q$. We define the function $f\left({y}\right):=y\log\frac{2y}{y+1}+\log\frac{2}{y+1}$. The Jensen-Shannon (JS) divergence is then defined as $D_{JS}\left({P||Q}\right)=\int_{\mathcal{X}}f\left({{dP\left({x}\right)}/{dQ\left({x}\right)}}\right)dQ\left({x}\right)$. Then we use a parametric based model ${w}_{\psi}\left({x}\right)$ to represent density ratio $\frac{dP}{dQ}$ and estimated the density ratio by maximizing the lower bound of $D_{JS}\left({P||Q}\right)$:
@@
3:Select critic EMA weight $\rho$, batch size $N$, determine number of Critic targets to subset $Z\in\left\{1,2\right\}$
@@
13:              Sample batch $b_{N}$ of size $N$ according to $P_{\mathcal{R}}$ from $\mathcal{R}_{N}$
@@
15:              With $b$, set $y=r+\gamma\Big{(}\min_{i\in\mathcal{Z}}Q_{\theta^{\prime}_{i}}\left({s^{\prime},a^{\prime}}\right)+\Big{.}$$\Big{.}\alpha\log\pi_{\phi}\left({a^{\prime}|s^{\prime}}\right)\Big{)},a^{\prime}\sim\pi_{\phi}\left({\cdot|s^{\prime}}\right)$
@@
20:         With b, update $\phi$ maximizing objective:
@@
Note that this advantage-aligned sampling strategy is not a heuristic-based approach but is theoretically derived in the performance difference lemma [22], providing insights into its effectiveness and superiority over the random sampling approach (see Section 5).
@@
5 Theoretical Analysis
In this section, we derive the priority term theoretically from the performance difference lemma [22] and show that our active advantage-aligned sampling strategy leads to improved policy performance. Furthermore, we establish a theoretical lower bound on the performance improvement gap under our sampling scheme.
@@ 5 Theoretical Analysis
Suppose the Q-function class is uniformly bounded, and for any Q-function, the corresponding optimal policy lies within the policy function class. Let $\epsilon^{t}$ denote the $\ell_{2}$ error of the Q-function in the critic update step. Let $\pi^{t}$ be the policy at iteration $t$ in $A^{3}$, updated using priority-weighted sampling with $w(s,a)\exp(\xi\cdot A(s,a))$. Then, the following lower bound holds:
@@
$R^{t}(s,a;\xi)$ $=\left(\frac{{\pi^{t+1}}(a{\,|\,}s)}{d^{\text{on}}(a{\,|\,}s)}\right)^{1-\xi}\cdot\frac{\sum_{s^{\prime},a^{\prime}}d^{\text{on}}(a^{\prime},s^{\prime})\pi^{t+1}(a^{\prime}{\,|\,}s^{\prime})^{\xi}}{d^{\text{on}}(a{\,|\,}s)^{\xi}}\cdot\frac{d^{\pi^{t+1}}(s)}{d^{\text{on}}(s)}.$
The proof is provided in Appendix A. We note that the coefficient $R^{t}(s,a;\xi)$ is not necessarily the tightest possible bound, since it is based on the supremum norm and therefore can be dominated by a single $(s,a)$ pair. A sharper result could be obtained by measuring distribution shift in the $\ell_{2}$ norm (or some other weaker norm). We nevertheless adopt the simpler supremum-norm bound here for clarity and to highlight the core intuition behind why advantage reweighting yields improvement, as will be detailed in the following.
@@
$\frac{d^{\pi^{t+1}}(s,a)}{\rho(s,a)}=$ $\left(\frac{{\pi^{t+1}}(a{\,|\,}s)}{d^{\text{on}}(a{\,|\,}s)}\right)^{1-\xi}\cdot\frac{\sum_{s^{\prime},a^{\prime}}d^{\text{on}}(a^{\prime},s^{\prime})\pi^{t+1}(a^{\prime}{\,|\,}s^{\prime})^{\xi}}{d^{\text{on}}(a{\,|\,}s)^{\xi}}\cdot\frac{d^{\pi^{t+1}}(s)}{d^{\text{on}}(s)},$
@@
a halfcheetah-random
b ant-random
c walker2d-medium
d antmaze-medium
Figure 1: Main results. A comparison between $A^{3}$, the state-of-the-art baseline RLPD and SAC with offline data (SACfD) on various D4RL benchmark tasks. (a-c): dense reward, (d): sparse reward. Shaded areas represent one standard deviation based on ten seeds.
6 Experiments
Environments
We evaluate $A^{3}$ on both dense and sparse reward tasks from the D4RL benchmark [12]. These include halfcheetah, walker2d, and ant, which are dense reward locomotion tasks, and antmaze, which involves sparse rewards. Each environment offers offline datasets composed of trajectories ranging from completely random to expert. Appendix C provides additional details.
Setup. We employ the basic setup of the SAC networks as recommended by [2], i.e., with an ensemble of size 10 each for critic networks and target critic networks, as well as entropy regularization. A significant difference is that the MLP underlying these networks only has 2 layers of size 256 each, as we desired to see if the agent is able to learn with less complexity.
Baseline Methods
For our main results, we compare $A^{3}$ with two baselines: (1) RLPD [2], regarded as the SOTA baseline for addressing online RL with offline datasets, also attains state-of-the-art performance in this problem set, (2) SAC with offline data (SACfD), a canonical off-policy approach using offline data, as also studied in [43] and [64]. In the ablation studies, we evaluate $A^{3}$ against five additional representative baselines: (3) Off2On [29], an offline-to-online RL method; (4) a variant of $A^{3}$ using advantage estimation only; (5) an online version of $A^{3}$ that excludes offline data; (6) SAC in an online setting without offline data; (7) TD (Temporal Difference) with a PER [55] sampling strategy; and (8) TD+Density, which combines PER with a density ratio sampling strategy.
6.1 Main results
Fig. 1 presents a comparative analysis of $A^{3}$’s performance against the baseline SACfD and the current state-of-the-art method, RLPD. The results demonstrate that $A^{3}$ consistently outperforms the baseline across the evaluated domains. This performance advantage can be attributed to a fundamental difference in sampling strategy: while RLPD relies on symmetric random sampling, $A^{3}$ employs an active sampling approach based on advantage alignment.
Unlike RLPD, which treats all transitions uniformly, $A^{3}$ dynamically reevaluates the relevance and on-policyness of each transition as the policy evolves, continuously adjusting its sampling priority to align with the current learning needs. This targeted sampling ensures that the most beneficial transitions are prioritized, directly contributing to faster and more effective policy improvement.
In scenarios involving nearly random offline datasets  Fig.1a,1b, datasets containing trajectories from a poorly performing policy, or even medium datasets Fig.1d,1c, useful transitions are often sparse and scattered. Random sampling, as used by RLPD, is likely to miss these valuable data points, leading to suboptimal performance. In contrast, $A^{3}$ ’s active sampling strategy effectively identifies and emphasizes these critical transitions, resulting in substantial policy enhancements, as clearly illustrated in Fig. 1.
Figure 2: halfcheetah-expert
In expert environments Fig. 2, $A^{3}$ consistently matches or exceeds the performance of RLPD. This superior performance can be attributed to the higher quality of transitions present in medium and expert datasets, compared to random datasets. Consequently, even with a random sampling strategy, RLPD is still likely to encounter useful transitions. However, most offline datasets are provided in a black-box format, where the specifics of the data are unknown. Despite this uncertainty, $A^{3}$ achieves performance that is at least on par with RLPD, demonstrating robustness to varying data quality in these black-box conditions.
6.2 Ablation studies
Ablation on advantage term
Fig. 3a illustrates the comparison between the performance of $A^{3}$ using advantage-aligned sampling priority and Off2On utilizing solely density ratio (${\sigma}=\mathbb{I}^{\text{off}}{w}\left({s_{i},a_{i}}\right)+\mathbb{I}^{\text{on}}$), a modified version of balanced experience replay [29]. The results show that $A^{3}$ with the advantage term surpasses its counterpart that only considers online-ness in prioritizing samples in the halfcheetah-random environments. This superiority is attributed to the advantage term, which effectively screens out transitions that are either non-informative or harmful. For example, even if a transition indicates online-ness, it may not provide new information if the policy has already mastered the associated action for that state. By integrating the advantage term, such repetitive transitions are excluded, as the advantage value tends to zero for well-understood transitions.
@@ 6.2 Ablation studies
Ablation on density term
@@
Fig. 3b compares the performance of $A^{3}$ to $A^{3}$ with only advantage in sampling priority (${\sigma}=\exp\left({\mathcal{\xi}\cdot A}\right)$), without density term. The results consistently show that $A^{3}$, which incorporates onlineness through the density term $w=d^{\text{on}}/d^{\text{off}}$, outperforms the version that does not. Onlineness measures the likelihood that $A^{3}$ will experience the given transition during the online exploration and exploitation of the current policy. Transitions experienced during online policy enhancement are more advantageous for policy development. In contrast, focusing on transitions that are unlikely to occur during live interactions with the environment can hinder the progression of policy improvement. This result demonstrates the effectiveness of onlineness term.
Ablation on purely online setting and offline data
@@
Fig. 3c presents an ablation study comparing regular $A^{3}$ (in red), purely online $A^{3}$ (in blue), and SAC (in green), with neither having access to offline data. $A^{3}$ surpasses its purely online version when utilizing an offline dataset, as the offline data provides a more diverse range of transitions that the online policy might not encounter, effectively demonstrating $A^{3}$’s ability to leverage offline datasets. Moreover, the purely online version of $A^{3}$ outperforms SAC, highlighting $A^{3}$’s robustness in environment setting. The results confirm $A^{3}$’s effectiveness in a purely online environment and its superiority over SAC in online batch scenarios through active advantage-aligned sampling.
Ablation on priority term
@@
Fig. 3d presents an ablation study for $A^{3}$ (in red), where we compare two different sampling strategies: PER as detailed in [55] (named as TD in blue), and a modified version incorporating a density ratio (named as TD+Density in green). The TD-error based sampling strategy prioritizes transitions with larger TD-errors. $A^{3}$ significantly outperforms both strategies, illustrating that an active advantage-aligned sampling approach is more effective than prioritizing based on TD-error alone. The superior performance of $A^{3}$ over TD+Density also indicates that prioritizing using the advantage term achieve the better performance compared to the TD-error term.
7 Conclusion
@@
We present $A^{3}$, a novel algorithm for online RL with offline dataset through a confidence-aware active advantage-aligned sampling strategy. This algorithm is theoretically motivated by the objective of shifting the sampling distribution toward more beneficial transitions to maximize policy improvement. We provide theoretical insights for $A^{3}$ and quantify its enhancement gap. Moreover, we conduct comprehensive experiments with various qualities of offline data, demonstrating that $A^{3}$ outperforms the SOTA RLPD method with significance. We also conduct multiple ablation studies and confirm the importance of each component within the active advantage-aligned formula and its effectiveness to pure online setting as well. While our approach primarily aims to enhance performance, it may result in higher computational costs due to the calculations needed for determining advantage-aligned sampling priorities. Reducing computational demands will be a focus of our future work.
@@ Appendix A Theoretical Motivation
In this section, we show that the active advantage-aligned sampling strategy helps mitigate the gap between offline data distribution, online data distribution and the current on-policy distribution, which serves as a main theoretical motivation for designing $A^{3}$.
@@
where ^$\widehat{}Q^{\widetilde{\pi}}$ is the estimated Q function at the current iteration. In the above formula, $\mu$ is the sampled data distribution and $\sigma$ is the quantity calculated in (8). Suppose we take some function class $\pi_{\phi}$ which contains the optimal one-step policy improvement $\pi^{\star}$ and also the optimization target $\widehat{\pi}$. Using a shift of distribution, we have
@@
$\geq-\sup_{s,a}\left|\frac{\pi^{\star}(a{\,|\,}s)}{\widehat{\pi}(a{\,|\,}s)}-1\right|\cdot\mathbb{E}_{d^{\widehat{\pi}}}[|Q^{\widetilde{\pi}}-\widehat{}Q^{\widetilde{\pi}}|]\geq-C\cdot\mathbb{E}_{d^{\widehat{\pi}}}[|Q^{\widetilde{\pi}}-\widehat{}Q^{\widetilde{\pi}}|]$
@@
$\mathbb{E}_{d^{\widehat{\pi}}}[|Q^{\widetilde{\pi}}-\widehat{}Q^{\widetilde{\pi}}|]$ $=\mathbb{E}_{\rho}\left[|Q^{\widetilde{\pi}}-\widehat{}Q^{\widetilde{\pi}}|\cdot\frac{d^{\widehat{\pi}}(s,a)}{\rho(s,a)}\right]\leq\sqrt{\mathbb{E}_{\rho}[(Q^{\widetilde{\pi}}-\widehat{}Q^{\widetilde{\pi}})^{2}]}\cdot\sup_{s,a}\left|\frac{d^{\widehat{\pi}}(s,a)}{\rho(s,a)}\right|.$
@@
Appendix B Additional Related Work
Offline to online RL

### v3

@@ Introduction
Estimating the likelihoods $d^{\text{off}}\left({s,a}\right)$ and $d^{\text{on}}\left({s,a}\right)$ poses a challenge, as they could represent stationary distributions from mixture of complex policy. To address this issue, we employ a method studied by Lee et al. (2022); Sinha et al. (2022) for density ratio estimation that does not rely on likelihoods. This method approximates ${w}\left({s,a}\right)$ by training a neural network ${w}_{\psi_{i}}\left({s,a}\right)$, which is parameterized by $\psi_{i},i\in[N_{e}]$, where $N_{e}$ is the number of density networks in the ensemble. The training exclusively uses samples from $\mathcal{B}^{\text{off}}$ and $\mathcal{B}^{\text{on}}$. We use variational representation of f-divergences (Nguyen, Wainwright, and Jordan 2007). Consider $P$ and $Q$ as probability measures on a measurable space $\mathcal{X}$, with $P$ being absolutely continuous $w.r.t$ $Q$. We define the function $f\left({y}\right):=y\log\frac{2y}{y+1}+\log\frac{2}{y+1}$. The Jensen-Shannon (JS) divergence is then defined as $D_{JS}\left({P||Q}\right)=\int_{\mathcal{X}}f\left({{dP\left({x}\right)}/{dQ\left({x}\right)}}\right)dQ\left({x}\right)$. Then we use a parametric based model ${w}_{\psi}\left({x}\right)$ to represent density ratio $\frac{dP}{dQ}$ and estimated the density ratio by maximizing the lower bound of $D_{JS}\left({P||Q}\right)$:
@@
3:Select critic EMA weight $\rho$, batch size NN, determine number of Critic targets to subset $Z\in\left\{1,2\right\}$
@@
13:     Sample batch $b_{N}$ of size NN according to $P_{\mathcal{R}}$ from $\mathcal{R}_{N}$
@@
15:     With $b_{N}$, set $y=r+\gamma\Big{(}\min_{i\in\mathcal{Z}}Q_{\theta^{\prime}_{i}}\left({s^{\prime},a^{\prime}}\right)+\Big{.}$ $\Big{.}\alpha\log\pi_{\phi}\left({a^{\prime}|s^{\prime}}\right)\Big{)},a^{\prime}\sim\pi_{\phi}\left({\cdot|s^{\prime}}\right)$
@@
20:   With $b_{N}$, update $\phi$ maximizing objective:
@@
Note that this advantage-aligned sampling strategy is not a heuristic-based approach but is theoretically derived in the performance difference lemma (Kakade and Langford 2002), providing insights into its effectiveness and superiority over the random sampling approach (see Section Theoretical Analysis ).
@@
Theoretical Analysis
In this section, we derive the priority term theoretically from the performance difference lemma (Kakade and Langford 2002) and show that our active advantage-aligned sampling strategy leads to improved policy performance. Furthermore, we establish a theoretical lower bound on the performance improvement gap under our sampling scheme.
@@
Suppose the Q-function class is uniformly bounded, and for any Q-function, the corresponding optimal policy lies within the policy function class. Let $\epsilon^{t}$ denote the $\ell_{2}$ error of the Q-function in the critic update step. Let $\pi^{t}$ be the policy at iteration $t$ in $A^{3}$RL, updated using priority-weighted sampling with $w(s,a)\exp(\xi\cdot A(s,a))$. Then, the following lower bound holds:
@@
$R^{t}(s,a;\xi)$ $=\left(\frac{{\pi^{t+1}}(a{\,|\,}s)}{d^{\text{on}}(a{\,|\,}s)}\right)^{1-\xi}$
$\cdot\frac{\sum_{s^{\prime},a^{\prime}}d^{\text{on}}(a^{\prime},s^{\prime})\pi^{t+1}(a^{\prime}{\,|\,}s^{\prime})^{\xi}}{d^{\text{on}}(a{\,|\,}s)^{\xi}}\cdot\frac{d^{\pi^{t+1}}(s)}{d^{\text{on}}(s)}.$
The proof is provided in the Appendix. We note that the coefficient $R^{t}(s,a;\xi)$ is not necessarily the tightest possible bound, since it is based on the supremum norm and therefore can be dominated by a single $(s,a)$ pair. A sharper result could be obtained by measuring distribution shift in the $\ell_{2}$ norm (or some other weaker norm). We nevertheless adopt the simpler supremum-norm bound here for clarity and to highlight the core intuition behind why advantage reweighting yields improvement, as will be detailed in the following.
a halfcheetah-random
b ant-random
c walker2d-medium
d antmaze-medium
Figure 2: Main results. A comparison between $A^{3}$RL, the state-of-the-art baseline RLPD and SAC with offline data (SACfD) on various D4RL benchmark tasks. (a-c): dense reward, (d): sparse reward. Shaded areas represent one standard deviation based on ten seeds.
@@
$\frac{d^{\pi^{t+1}}(s,a)}{\rho(s,a)}=$ $\left(\frac{{\pi^{t+1}}(a{\,|\,}s)}{d^{\text{on}}(a{\,|\,}s)}\right)^{1-\xi}$
$\cdot\frac{\sum_{s^{\prime},a^{\prime}}d^{\text{on}}(a^{\prime},s^{\prime})\pi^{t+1}(a^{\prime}{\,|\,}s^{\prime})^{\xi}}{d^{\text{on}}(a{\,|\,}s)^{\xi}}\cdot\frac{d^{\pi^{t+1}}(s)}{d^{\text{on}}(s)},$
@@
Experiments
@@
Figure 4: halfcheetah-expert
Environments.
We evaluate $A^{3}$RL on both dense and sparse reward tasks from the D4RL benchmark (Fu et al. 2020). These include halfcheetah, walker2d, and ant, which are dense reward locomotion tasks, and antmaze, which involves sparse rewards. Each environment offers offline datasets composed of trajectories ranging from completely random to expert. We defer additional details on the environment to the Appendix.
Setup.
We employ the basic setup of the SAC networks as recommended by (Ball et al. 2023), i.e., with an ensemble of size 10 each for critic networks and target critic networks, as well as entropy regularization. A significant difference is that the MLP underlying these networks only has 2 layers of size 256 each, as we desired to see if the agent is able to learn with less complexity.
Baseline Methods.
For our main results, we compare $A^{3}$RL with two baselines: (1) RLPD (Ball et al. 2023), regarded as the SOTA baseline for addressing online RL with offline datasets, also attains state-of-the-art performance in this problem set, (2) SAC with offline data (SACfD), a canonical off-policy approach using offline data, as also studied in (Nair et al. 2020) and (Vecerik et al. 2017). In the ablation studies, we evaluate $A^{3}$RL against five additional representative baselines: (3) Off2On (Lee et al. 2022), an offline-to-online RL method; (4) a variant of $A^{3}$RL using advantage estimation only; (5) an online version of $A^{3}$RL that excludes offline data; (6) SAC in an online setting without offline data; (7) TD (Temporal Difference) with a PER (Schaul et al. 2015) sampling strategy; and (8) TD+Density, which combines PER with a density ratio sampling strategy.
Main results
Fig. 2 presents a comparative analysis of $A^{3}$RL’s performance against the baseline SACfD and the current state-of-the-art method, RLPD. The results demonstrate that $A^{3}$RL consistently outperforms the baseline across the evaluated domains. This performance advantage can be attributed to a fundamental difference in sampling strategy: while RLPD relies on symmetric random sampling, $A^{3}$RL employs an active sampling approach based on advantage alignment.
Unlike RLPD, which treats all transitions uniformly, $A^{3}$RL dynamically reevaluates the relevance and on-policyness of each transition as the policy evolves, continuously adjusting its sampling priority to align with the current learning needs. This targeted sampling ensures that the most beneficial transitions are prioritized, directly contributing to faster and more effective policy improvement.
In scenarios involving nearly random offline datasets  Fig.2a,2b, datasets containing trajectories from a poorly performing policy, or even medium datasets Fig.2d,2c, useful transitions are often sparse and scattered. Random sampling, as used by RLPD, is likely to miss these valuable data points, leading to suboptimal performance. In contrast, $A^{3}$RL ’s active sampling strategy effectively identifies and emphasizes these critical transitions, resulting in substantial policy enhancements, as clearly illustrated in Fig. 2.
In expert environments Fig. 4, RLPD performs on par with $A^{3}$RL. This superior performance can be attributed to the higher quality of transitions present in medium and expert datasets, compared to random datasets. Consequently, even with a random sampling strategy, RLPD is still likely to encounter useful transitions. However, most offline datasets are provided in a black-box format, where the specifics of the data are unknown. Despite this uncertainty, $A^{3}$RL achieves performance that is at least on par with RLPD, demonstrating robustness to varying data quality in these black-box conditions.
Ablation studies
Ablation on advantage term.
Fig. 3a illustrates the comparison between the performance of $A^{3}$RL using advantage-aligned sampling priority and Off2On utilizing solely density ratio (${\sigma}=\mathbb{I}^{\text{off}}{w}\left({s_{i},a_{i}}\right)+\mathbb{I}^{\text{on}}$), a modified version of balanced experience replay (Lee et al. 2022). The results show that $A^{3}$RL with the advantage term surpasses its counterpart that only considers online-ness in prioritizing samples in the halfcheetah-random environments. This superiority is attributed to the advantage term, which effectively screens out transitions that are either non-informative or harmful. For example, even if a transition indicates online-ness, it may not provide new information if the policy has already mastered the associated action for that state. By integrating the advantage term, such repetitive transitions are excluded, as the advantage value tends to zero for well-understood transitions.
Ablation on density term.
@@
Fig. 3b compares the performance of $A^{3}$RL to $A^{3}$RL with only advantage in sampling priority (${\sigma}=\exp\left({\mathcal{\xi}\cdot A}\right)$), without density term. The results consistently show that $A^{3}$RL, which incorporates onlineness through the density term $w=d^{\text{on}}/d^{\text{off}}$, outperforms the version that does not. Onlineness measures the likelihood that $A^{3}$RL will experience the given transition during the online exploration and exploitation of the current policy. Transitions experienced during online policy enhancement are more advantageous for policy development. In contrast, focusing on transitions that are unlikely to occur during live interactions with the environment can hinder the progression of policy improvement. This result demonstrates the effectiveness of onlineness term.
Ablation on purely online setting and offline data.
@@
Fig. 3c presents an ablation study comparing regular $A^{3}$RL (in red), purely online $A^{3}$RL (in blue), and SAC (in green), with neither having access to offline data. $A^{3}$RL surpasses its purely online version when utilizing an offline dataset, as the offline data provides a more diverse range of transitions that the online policy might not encounter, effectively demonstrating $A^{3}$RL’s ability to leverage offline datasets. Moreover, the purely online version of $A^{3}$RL outperforms SAC, highlighting $A^{3}$RL’s robustness in environment setting. The results confirm $A^{3}$RL’s effectiveness in a purely online environment and its superiority over SAC in online batch scenarios through active advantage-aligned sampling.
Ablation on priority term.
@@
Fig. 3d presents an ablation study for $A^{3}$RL (in red), where we compare two different sampling strategies: PER as detailed in (Schaul et al. 2015) (named as TD in blue), and a modified version incorporating a density ratio (named as TD+Density in green). The TD-error based sampling strategy prioritizes transitions with larger TD-errors. $A^{3}$RL significantly outperforms both strategies, illustrating that an active advantage-aligned sampling approach is more effective than prioritizing based on TD-error alone. The superior performance of $A^{3}$RL over TD+Density also indicates that prioritizing using the advantage term achieve the better performance compared to the TD-error term.
Conclusion
@@
We present $A^{3}$RL, a novel algorithm for online RL with offline dataset through a confidence-aware active advantage-aligned sampling strategy. This algorithm is theoretically motivated by the objective of shifting the sampling distribution toward more beneficial transitions to maximize policy improvement. We provide theoretical insights for $A^{3}$RL and quantify its enhancement gap. Moreover, we conduct comprehensive experiments with various qualities of offline data, demonstrating that $A^{3}$RL outperforms the SOTA RLPD method with significance. We also conduct multiple ablation studies and confirm the importance of each component within the active advantage-aligned formula and its effectiveness to pure online setting as well. While our approach primarily aims to enhance performance, it may result in higher computational costs due to the calculations needed for determining advantage-aligned sampling priorities. Reducing computational demands will be a focus of our future work.
@@ Appendix A Theoretical Motivation
In this section, we show that the active advantage-aligned sampling strategy helps mitigate the gap between offline data distribution, online data distribution and the current on-policy distribution, which serves as a main theoretical motivation for designing $A^{3}$RL.
@@
where $\widehat{}Q^{\widetilde{\pi}}$ is the estimated Q function at the current iteration. In the above formula, $\mu$ is the sampled data distribution and $\sigma$ is the quantity calculated in (9). Suppose we take some function class $\pi_{\phi}$ which contains the optimal one-step policy improvement $\pi^{\star}$ and also the optimization target $\widehat{\pi}$. Using a shift of distribution, we have
@@
[…]
@@
[…]
@@
Appendix B Additional Preliminaries
Layer Normalization: Off-policy RL algorithms often query the learned QQ–function with out-of-distribution actions, leading to overestimation errors due to function approximation. This can cause training instabilities and even divergence, particularly when the critic struggles to keep up with growing value estimates. To address this, prior research has employed Layer Normalization to ensure that the acquired functions do not extrapolate in an unconstrained manner. Layer Normalization acts to confine $Q$-values within the boundaries set by the norm of the weight layer, even for actions beyond the dataset. As a result, the impact of inaccurately extrapolated actions is substantially reduced, as their associated $Q$-values are unlikely to significantly exceed those already observed in the existing data. Consequently, Layer Normalization serves to alleviate issues such as critic divergence and the occurrence of catastrophic overestimation.
Maximum Entropy RL: Incorporating entropy into the learning objective (as defined in (1)) helps mitigate overconfidence in value estimates, particularly when training with offline datasets. In offline RL, policies may become overly conservative due to limited dataset coverage, leading to suboptimal exploration during fine-tuning. By preserving policy stochasticity, entropy regularization ensures that the agent remains adaptable when transitioning from offline training to online interactions. This controlled exploration has been shown to improve training stability and prevent premature convergence (Ball et al. 2023; Chen et al. 2021; Haarnoja et al. 2018; Hiraoka et al. 2021).
Appendix C Additional Related Work
Offline to online RL.